Find the monthly payment that will yield the given future value. (Round your answer to the nearest cent.)

$105,000 at 9 1/4

% interest for thirty-two years; ordinary annuity
Answer #1

Answer is $44.77

FIND PMT

Rate=9.25%/12

NPER=32*12=384

FV=$105000

PV=0

=PMT(9.25%/12,384,0,-105000,0)=$44.77